OpenStudy (anonymous):

whats the answer for ^x-1=-3 1. 9 2. 16 3. 10 4. no solution

OpenStudy (anonymous):

Retype your question...

OpenStudy (anonymous):

the square root symbol over x-1 = -3

OpenStudy (saifoo.khan):

draw it.

OpenStudy (anonymous):



OpenStudy (sgadi):

3. x=10 \[\sqrt{x-1}=-3\] squaring on both sides, we get \[\Rightarrow {x-1}=9\] \[\Rightarrow {x}=9+1=10\]

jhonyy9 (jhonyy9):

3. is right